OpenStudy (anonymous):

Triangles DEF and HJK are similar. DEF has side lengths DE = 10, EF = 15, and FD = 20. Which of the following could be the side lengths of HJK? A.8, 10, 12 B.5, 7, 9 C.2, 3, 4 D.2, 6, 7

OpenStudy (anonymous):

{2,3,4} \[ \frac{10}2 =\frac{15}3=\frac {20}{4}=5 \]

OpenStudy (anonymous):

So..then its... c?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

yw